### v2

@@ Abstract
Causal weighted quantile treatment effects (WQTE) are a useful complement to standard causal contrasts that focus on the mean when interest lies at the tails of the counterfactual distribution. To-date, however, methods for estimation and inference regarding causal WQTEs have assumed complete data on all relevant factors. In most practical settings, however, data will be missing or incomplete data, particularly when the data are not collected for research purposes, as is the case for electronic health records and disease registries. Furthermore, such data sources may be particularly susceptible to the outcome data being missing-not-at-random (MNAR). In this paper, we consider the use of double-sampling, through which the otherwise missing data are ascertained on a sub-sample of study units, as a strategy to mitigate bias due to MNAR data in the estimation of causal WQTEs. With the additional data in-hand, we present identifying conditions that do not require assumptions regarding missingness in the original data. We then propose a novel inverse-probability weighted estimator and derive its asymptotic properties, both pointwise at specific quantiles and uniformly across a range of quantiles over some compact subset of $(0,1)$, allowing the propensity score and double-sampling probabilities to be estimated. For practical inference, we develop a bootstrap method that can be used for both pointwise and uniform inference. A simulation study is conducted to examine the finite sample performance of the proposed estimators. The proposed method is illustrated with data from an EHR-based study examining the relative effects of two bariatric surgery procedures on BMI loss at 3 years post-surgery.
@@
Bootstrap, Causal inference, Quantile regression, Quantile regression process, Heterogeneous treatment effect, Missing not at random, Uniform inference, Unspecified missing data mechanism.
@@ 1 Introduction
Modern causal inference methods conceptualize causation via a comparison of counterfactuals or potential outcomes across interventions or treatment strategies (Hernán and Robins, 2020). Because one cannot hope to learn about causation for individuals, it is standard to focus on contrasts of some summary statistic of the treatment-specific counterfactual distributions. The most common such contrast, arguably, is the average treatment effect (ATE), although related contrasts have been considered including the average treatment effect on the treated (e.g., Rubin, 1977; Imbens, 2004), the survivor average treatment effect (e.g., Robins, 1986; Tchetgen Tchetgen, 2014), and average effects of modified treatment policies (e.g., Muñoz and van der Laan, 2012; Haneuse and Rotnitzky, 2013). That each of these contrasts focuses on the mean of the counterfactual distribution, in one way or another, is natural since the mean is standard measure of centrality. In some settings, however, particularly when low-probability but high-impact events/outcomes are anticipated, understanding the causal effect of a given treatment beyond the ‘center’, including at the tails of the counterfactual distributions, may be more relevant to decision-making (Sun et al., 2021; Meyerson et al., 2023). In such settings, quantile treatment effects (QTEs), defined as the difference in quantiles of counterfactual distributions across treatment options, may be a useful compliment to standard contrasts based on the mean (e.g., Andrews et al., 2016; Schiele and Schmitz, 2016; Autor et al., 2017; Hemilä et al., 2022).
@@
Towards estimation of causal QTEs, there is an emerging literature most of which focuses on inverse-probability weighting (IPW) using the propensity score (Rosenbaum and Rubin, 1983; Firpo, 2007; Sun et al., 2021). That methods analogous to the g-formula or doubly-robust methods have not been developed is, in part, due to Jensen’s inequality precluding the characterization of population QTEs as a function of covariate-conditional QTEs. Throughout this literature, and in contrast to the literature on related but distinct quantile regression methods (Yi and He, 2009; Yuan and Yin, 2010; Wei et al., 2012; Zhao et al., 2013; Sherwood et al., 2013; Han et al., 2019; Yu et al., 2023), it has been implicitly assumed that complete data is available on the outcome and all relevant covariates. Missing or incomplete data, however, is a well-known and widespread practical challenge in almost all research settings where human subjects are involved. This is particularly the case when the data at-hand were not collected for research purposes as in the case with electronic health records (EHRs), claims-based databases and registries (Haneuse and Shortreed, 2017). In studies of the long-term effects of bariatric surgery, for example, data on relevant outcomes (e.g., weight/BMI, hypertension, type 2 diabetes and mortality) are subject to a range of mechanisms that result in missing data, including: disenrollment from a health plan; that a patient doesn’t have a relevant health encounter at a time point of interest; and, that a measurement was not taken even though standards of care would indicate that it should have been (Haneuse and Daniels, 2016; Peskoe et al., 2021).
@@
When it comes to performing analyses based on data that are incomplete, however, analysts must engage in two related tasks: (i) clarifying and invoking assumptions necessary for valid estimation and inference; and, (ii) employing some statistical methodology that acknowledges the missingness, as appropriate. Regarding the former, it is typical that the missing-at-random (MAR) assumption is invoked (Little and Rubin, 2019); intuitively, MAR corresponds to probability of missingness depending only on observed data. In some settings, however, MAR may not hold, in that missingness depends, in part at least, on quantities that are unobserved. This scenario is generally referred to as missing-not-at-random (MNAR) or informative missingness. Unfortunately, in the absence of additional information, MNAR cannot be empirically verified. In turn, much of the literature has focused on sensitivity analysis methods (e.g., Robins et al., 2000) and estimating bounds on the parameters of interest (e.g., Manski, 1990). An alternative approach, which we pursue in this work, is to engage in additional data collection efforts, specifically to (re)contact a subset of subjects with missing data to obtain complete data. With this data in-hand, otherwise untestable assumptions can be evaluated (at least partially) and/or incorporated into analyses. Such a strategy, sometimes referred to as double-sampling is well-established as a means to mitigate confounding bias (Breslow, 1996; Schildcrout et al., 2020), bias due to measurement error (Lynn et al., 2012; Bound et al., 2001) and ecological bias (Haneuse and Wakefield, 2007, 2008). However, with few exceptions (Frangakis and Rubin, 2001; An et al., 2009; Qian et al., 2020; Coppock et al., 2017; Levis et al., 2023), double-sampling has not been considered routinely as a means to mitigate bias due to potentially MNAR data.
@@
With this collective backdrop, this paper proposes the use of double-sampling as a means to mitigate bias due to potentially informative missing data for a continuous outcome when interest lies in estimating contrasts in a class of weighted quantile treatment effects (WQTEs). These estimands, introduced by Sun et al. (2021), facilitate comparisons of potential outcomes over various selected target populations, such as among the treated. Towards this, we first put forth conditions sufficient for the identification of WQTEs in observational settings with missing outcomes, when the double-sampling design is pursued. Based on this, we propose an IPW approach to estimate WQTEs, labelled IPW-WQTE. We derive the asymptotic distribution of the proposed estimator with estimated propensity scores and missing mechanism at any fixed quantile level (i.e., pointwise inference), and establish uniform inference of the entire IPW-WQTEs process across all quantiles over a compact subset of $(0,1)$. Finally, we propose a bootstrap resampling-based method can be used for either pointwise or uniform inference. The performance of the proposed approach and estimator is evaluated via a simulation study, with proofs of the main theorems provided in the Supplementary Material. The proposed method is then illustrated with data from an EHR-based study examining the relative effects of two bariatric surgery procedures on BMI loss at 3 years post-surgery.
@@ 2.1 Notation
Let $Y\in\mathbbm{R}$ denote the continuous outcome of interest and $Z\in\mathcal{Z}=\{0,1\}$ a binary treatment indicator (e.g., $Z$=$0/1$=control/treatment). The methods proposed in this paper are derived under binary treatment for simplicity, but can be generalized to categorical and continuous treatment using the generalized propensity score (e.g., Imbens, 2000; Sun et al., 2021). Let $X\in\mathcal{X}\subset\mathbbm{R}^{p}$ denote a $p$-dimensional vector of covariates that are sufficient for the control of confounding of the $Z\sim Y$ association. We refer to $(Y,Z,X^{\top})$ as the ‘complete data’ and conceive of it as arising from some joint distribution $P_{c}$ from which one could, in principle, derive, among other quantities: $F_{Y}(y)=\textrm{pr}(Y\leq y)$, the marginal cumulative distribution function of $Y$; and, $F_{Y\mid Z,X}(y)$ and $f_{Y\mid Z,X}(y)$ the conditional cumulative distribution and density functions of $Y$ given $Z$ and $X$, respectively.
Towards embedding this work within the causal inference paradigm, let $Y(z)$ denote the potential outcome corresponding to treatment level $Z=z$ for $z\in\mathcal{Z}$, and $F_{Y(z)}(y)$ and $F_{Y(z)\mid X}(y)$ denote the corresponding marginal and conditional cumulative distribution functions, respectively.
@@ 2.2 Weighted quantile treatment effects
where $Q^{g}_{Y(z)}(\cdot)$ is the quantile function of $F^{g}_{Y(z)}(y)$. Note, the introduction of $g(\cdot)$ permits the specification of alternative target populations to which the intent is to generalize. For example, the population QTE (Koenker and Bassett Jr, 1978; Koenker, 2005) is a special case of the WQTE for $g(x)=1$. As another example, if $g(x)$ equals the propensity score $e(x)=\textrm{pr}(Z=1\mid X=x)$, the estimand in equation (2) corresponds to the $\tau$th QTE among the treated, a QTE analogue to the average treatment effect among the treated. See Sun et al. (2021) for more details.
@@ 2.3 Identification and estimation given complete data
$0=E\left[g(X)\frac{\mathbbm{1}\left\{Z=z\right\}}{e(X)^{z}(1-e(X))^{1-z}}\left(\tau-\mathbbm{1}\{Y\leq Q_{Y(z)}^{g}(\tau)\}\right)\right].$
These estimating equations motivate the following estimator: given a sample of independent and identically distributed (i.i.d.) draws from the joint distribution $P_{c}$, say $\{(Y_{1},Z_{1},X_{1}),\ldots,(Y_{n},Z_{n},X_{n})\}$, Sun et al. (2021) proposed the IPW Z-estimator defined as the solution to:
$\mathbbm{P}_{n}\left(W^{g}(Z,X)(1,Z)\left[\tau-\mathbbm{1}\{Y<\beta^{g}_{0}(\tau)+\beta^{g}(\tau)Z\}\right]\right)=0,\quad\tau\in(0,1),$ (3)
with vector $(1,Z)$ and
$W^{g}(Z_{i},X_{i})=\frac{g(X_{i})}{e(X_{i})}Z_{i}+\frac{g(X_{i})}{1-e(X_{i})}(1-Z_{i}).$ (4)
and where $\mathbbm{P}_{n}(U)=n^{-1}\sum_{i=1}^{n}U_{i}$ denotes the empirical mean of a generic random variable $U$.
@@ 3.1 Double-sampling
For settings where $Y$ is potentially missing, let $R\in\{0,1\}$ denote the corresponding observance indicator so that $R=0/1$ if $Y$ is missing/observed. In this context, MAR states, intuitively, that the probability of $Y$ being observed only depends on the fully observed variables. Formally, one can write MAR as a conditional independence assumption: $R\perp Y\mid Z,X$. From this formal definition, MNAR may be seen to result when missingness depends on the underlying value of the outcome, either directly or through some (as-yet unspecified) other factor(s). In the bariatric surgery context, for example, the missingness process may be MNAR if patients who do not respond to surgery (i.e., fail to lose weight and/or maintain weight loss) are systematically more likely to have complete data—even after accounting for observed covariates and treatment status—perhaps because they are more likely to continue engaging with the health care system.
In settings where it is not clear whether MAR holds, we consider collecting the otherwise missing information on a sub-sample of patients, a process termed ‘double-sampling’ by Frangakis and Rubin (2001). Towards this, let $S\in\{0,1\}$ be an indicator for double-sampling, specifically indicating whether an individual with initially missing outcome data is selected into the follow-up sample and complete data are successfully collected. Note, by design, $S$ can only be $1$ if $R=0$ and is deterministically 00 if $R=1$; therefore, $S\equiv S(1-R)$. Additionally, the process by which these otherwise missing data are ascertained will be context specific and may involve active participation of study subjects (e.g., response to a survey or phone call) or no participation (e.g., manual chart review); see Levis et al. (2023) for additional details on practicalities. Finally, under this scheme, the ‘observed data’ are $n$ i.i.d. copies of $\{R,S,(R+S)Y,Z,X^{\top}\}$ with corresponding joint distribution $P_{o}$.
@@ 3.2 Assumptions
At the outset, we present a series of assumptions that will be leveraged to establish identification of the effects of interest, and the asymptotic properties of the estimator proposed in Section 3.3:
Assumption 1.
@@
Assumption 2.
Positivity of double-sampling probabilities: For some $c_{\eta}>0$, $c_{\eta}<\eta(z,x)\coloneqq\textrm{pr}(S=1\mid Z=z,X=x,R=0)<1-c_{\eta}$ for $z\in\mathcal{Z},x\in\mathcal{X}$.
Assumption 3.
For all $\tau\in(0,1)$, $z\in\mathcal{Z}$, at the true values $(\beta^{g*}_{0}(\tau),\beta^{g*}(\tau))$, the distribution function $F_{Y\mid Z=z,X}$ has one-sided derivatives at $\beta^{g*}_{0}(\tau)+\beta^{g*}(\tau)z$ with density $f_{Y\mid Z=z,X}$.
Assumption 4.
Letting $(\beta^{g*}_{0}(\tau),\beta^{g*}(\tau))$ denote the true values of $(\beta^{g}_{0}(\tau),\beta^{g}(\tau))$ for all $\tau\in(0,1)$, we have
  1. 1.
    $D_{1}(\tau)=E\left[g(X)f_{Y(1)\mid X}\{\beta^{g*}_{0}(\tau)+\beta^{g*}(\tau)\}\right]\in(0,\infty)$;
  2. 2.
    $D_{0}(\tau)=E\left[g(X)f_{Y(0)\mid X}\{\beta^{g*}_{0}(\tau)\}\right]\in(0,\infty)$.
Assumption 5.
@@
Assumptions 1 and 2 relate directly to the double-sampling scheme and are required to ensure that the full data distribution, $P_{f}$, is identified given the observed data. Specifically, Assumption 1 states that whether or not an individual is double-sampled is independent of the initially unobserved outcome, conditional on initially observed data. Meanwhile, Assumption 2 requires that among individuals with initially missing data, the conditional probability of being double-sampled cannot be 0 or 1 in any subgroup defined by $(Z,X)$.
@@
Note, Assumption 1 can be viewed as an MAR assumption for the second phase of data collection. In practice, depending on the second phase sampling method, there may be a discrepancy between the intended sampling probabilities that will typically be under the control of the researchers, and the process by which data are actually collected (e.g., does an individual fill out the survey?). The follow-up sample, therefore, may or may not have complete data relative to the intended second stage sampling scheme. For example, when conducting a chart review without individuals’ direct engagement, it is reasonable to anticipate that those sub-sampled will provide complete data. In this setting, the plausibility of Assumption 1 may be quite likely. In other settings, however, such as when the data are collected via a telephone survey, some individuals initially selected by the design may decline to participate, leading to an additional layer of non-response (Koffman et al., 2021). In this situations, the double-sampling probabilities $\eta(Z,X)$ will have to be estimated, and the plausibility of Assumption 1 considered in light of this additional missingness.
Assumptions 3 and 4 are required to control the asymptotic behaviour of the proposed estimator. Briefly, Assumption 3 is standard in quantile-based analyses (Koenker and Bassett Jr, 1978; Koenker, 2005). Assumption 3 is necessary to ensure the expectation of the vector-valued estimating function ($\phi_{\theta}$ defined in expression (8)) is differentiable at the true values $\beta^{g*}_{0}(\tau)$ and $\beta^{g*}_{0}(\tau)+\beta^{g*}(\tau)$. Assumption 4 requires that the conditional density function be locally bounded such that the variance is finite, necessary to establish the asymptotic theorems in Section 3.4 and ‘Asymptotic theory for the entire quantile process’. We note that Assumption 4 is less restrictive than requiring the density function $f_{Y\mid Z,X}$ being continuous and uniformly bounded away from 0 and $\infty$ at $\beta^{g*}_{0}(\tau)+\beta^{g*}(\tau)Z_{i}$.
@@
Finally, Assumption 5 clarifies assumptions regarding parametric specifications for the double-sampling probabilities and propensity scores when one or both require estimation. Specifically, implicit to the assumption is that the class of models adopted for each is sufficiently broad as to contain the true underlying mechanism. Practically, consideration of this assumption will be made in the context of whether and how flexibility is introduced into models specifications (e.g., with the use of splines). While, arguably, a focus on parametric models may be restrictive, it likely aligns with how most related clinical and public health research is conducted, and so represents an important case to consider. Nevertheless, possible extensions to nonparametric models are discussed Section 6.

### v3

@@ Abstract
Causal weighted quantile treatment effects (WQTEs) complement standard mean-focused causal contrasts when interest lies at the tails of the counterfactual distribution. However, existing methods for estimating and inferring causal WQTEs assume complete data on all relevant factors, which is often not the case in practice, particularly when the data are not collected for research purposes, such as electronic health records (EHRs) and disease registries. Furthermore, these data may be particularly susceptible to the outcome data being missing-not-at-random (MNAR). This paper proposes to use double-sampling, through which the otherwise missing data are ascertained on a sub-sample of study units, as a strategy to mitigate bias due to MNAR data in estimating causal WQTEs. With the additional data, we present identifying conditions that do not require missingness assumptions in the original data. We then propose a novel inverse-probability weighted estimator and derive its asymptotic properties, both pointwise at specific quantiles and uniformly across quantiles over some compact subset of $(0,1)$, allowing the propensity score and double-sampling probabilities to be estimated. For practical inference, we develop a bootstrap method that can be used for both pointwise and uniform inference. A simulation study is conducted to examine the finite sample performance of the proposed estimators. We illustrate the proposed method using EHR data examining the relative effects of two bariatric surgery procedures on BMI loss three years post-surgery.
@@
Bootstrap, Causal inference, Heterogeneous treatment effect, Missing not at random, Quantile regression process, Uniform inference.
volume: 64
@@ 1 Introduction
Modern causal inference conceptualizes causation via a comparison of counterfactuals across interventions or treatment strategies (Hernán and Robins, 2020). As causation for individuals cannot be directly observed, analysis typically focuses on contrasts of some summary statistic of the treatment-specific counterfactual distributions. The most common contrast, arguably, is the average treatment effect (ATE), although related contrasts have been considered including the ATE on the treated (e.g., Imbens, 2004), the survivor ATE (e.g., Tchetgen Tchetgen, 2014), and average effects of modified treatment policies (e.g., Muñoz and van der Laan, 2012; Haneuse and Rotnitzky, 2013). That each of these contrasts focuses on the mean of the counterfactual distribution, in one way or another, is natural since the mean is standard measure of centrality. In some settings, however, particularly when low-probability but high-impact outcomes are anticipated, understanding the causal effect of a given treatment beyond the ‘center’, including at the tails of the counterfactual distributions, may be more relevant to decision-making (Sun et al., 2021; Meyerson et al., 2023). In such settings, quantile treatment effects (QTEs), defined as the difference in quantiles of counterfactual distributions across treatment options, may be a useful compliment to standard contrasts based on the mean (e.g., Andrews et al., 2016; Autor et al., 2017; Hemilä et al., 2022).
@@
Towards estimation of causal QTEs, there is an emerging literature most of which focuses on inverse-probability weighting (IPW) using the propensity score (Firpo, 2007; Sun et al., 2021). Throughout this literature, and in contrast to the literature on related but distinct quantile regression methods (Yuan and Yin, 2010; Sherwood et al., 2013; Han et al., 2019; Yu et al., 2023), it has been implicitly assumed complete data are available the outcome and all relevant covariates. Missing or incomplete data, however, is a well-known and widespread practical challenge in almost all research settings involving humans. This is particularly the case when the data at-hand were not collected for research purposes as in the case with electronic health records (EHRs), claims-based databases and registries (Haneuse and Shortreed, 2017). In studies of the long-term effects of bariatric surgery, for example, data on relevant outcomes (e.g., weight/BMI, hypertension, type 2 diabetes and mortality) are subject to a range of mechanisms that result in missing data, including: disenrollment from a health plan; that a patient does not have a relevant health encounter at a time point of interest; and, that a measurement was not taken even though standards of care would indicate that it should have been (Haneuse and Daniels, 2016; Peskoe et al., 2021).
@@
When it comes to perform analyzes based on data that are incomplete, however, analysts must engage in two related tasks: (i) clarifying and invoking assumptions necessary for valid estimation and inference; and, (ii) employing some statistical methodology that acknowledges the missingness, as appropriate. Regarding the former, it is typical that the missing-at-random (MAR) assumption is invoked (Little and Rubin, 2020); intuitively, MAR corresponds to probability of missingness depending only on observed data. In some settings, however, MAR may not hold, in that missingness depends, in part at least, on quantities that are unobserved. This scenario is generally referred to as missing-not-at-random (MNAR) or informative missingness. Unfortunately, in the absence of additional information, MNAR cannot be empirically verified. In turn, much of the literature has focused on sensitivity analysis methods (e.g., Robins et al., 2000) and estimating bounds on the parameters of interest (e.g., Manski, 1990). An alternative approach, which we pursue in this work, is to engage in additional data collection efforts, specifically to (re)contact a subset of subjects with missing data to obtain complete data. With this data in-hand, otherwise untestable assumptions can be evaluated (at least partially) and/or incorporated into analyzes. Such a strategy, sometimes referred to as double-sampling is well-established as a means to mitigate confounding bias (Breslow, 1996; Schildcrout et al., 2020), bias due to measurement error (Lynn et al., 2012; Bound et al., 2001) and ecological bias (Haneuse and Wakefield, 2007, 2008). However, with few exceptions (Frangakis and Rubin, 2001; An et al., 2009; Qian et al., 2020; Coppock et al., 2017; Levis et al., 2023), double-sampling has not been considered routinely as a means to mitigate bias due to potentially MNAR data.
@@
This paper proposes using double-sampling to mitigate bias due to potentially informative missing data for a continuous outcome when interest lies in estimating contrasts in a class of weighted quantile treatment effects (WQTEs). These estimands, introduced by Sun et al. (2021), facilitate comparisons of potential outcomes over various selected target populations, such as among the treated. We first put forth conditions sufficient for identifying WQTEs in observational settings with missing outcomes, when the double-sampling design is pursued. Based on this, we propose an IPW approach to estimate WQTEs, labelled IPW-WQTE. We derive the asymptotic distribution of the proposed estimator with estimated propensity scores and missing mechanism at any fixed quantile (pointwise inference), and establish uniform inference of the entire IPW-WQTEs process over a compact subset of $(0,1)$. Finally, we propose a bootstrap resampling method that can be used for either pointwise or uniform inference. The proofs of the main theorems provided in the Supplementary Material. The performance of the proposed approach and estimator is evaluated via a simulation study. The proposed method is then illustrated with data from an EHR-based study examining the relative effects of two bariatric surgery procedures on BMI loss at 3 years post-surgery.
@@ 2.1 Notation
Let $Y\in$ denote the continuous outcome of interest and $Z\in\mathcal{Z}=\{0,1\}$ a binary treatment indicator (e.g., $0/1$=control/treatment). The methods in this paper are derived under binary treatment for simplicity, but can be generalized to categorical and continuous treatment using the generalized propensity score (e.g., Imbens, 2000; Sun et al., 2021). Let $X\in\mathcal{X}\subset\mathbbm{R}^{p}$ denote a $p$-dimensional covariates vector sufficient to control for confounding. We refer to $(Y,Z,X)$ as the ‘complete data’ from a joint distribution $P_{c}$ used to derive the marginal cumulative distribution function $F_{Y}(y)=\textrm{pr}(Y\leq y)$, the conditional cumulative distribution $F_{Y\mid Z,X}(y)$, and density function $f_{Y\mid Z,X}(y)$. Let $\mathcal{Y}_{z,x}$ denote the support of the conditional distribution of $Y$ given $Z=z$ and $X=x$. Let $Y(z)$ denote the potential outcome under treatment $Z=z$, and $F_{Y(z)}(y)$ and $F_{Y(z)\mid X}(y)$ denote the corresponding marginal and conditional cumulative distribution functions, respectively, with densities $f_{Y(z)}(y)$ and $f_{Y(z)\mid X}(y)$. Let $\mathbbm{P}_{n}(U)=n^{-1}\sum_{i=1}^{n}U_{i}$ denote the empirical mean of a generic random variable $U$.
@@ 2.2 Weighted quantile treatment effects
where $Q^{g}_{Y(z)}(\cdot)$ is the quantile function of $F^{g}_{Y(z)}(y)$. Note, the introduction of $g(\cdot)$ permits the specification of alternative target populations to which the intent is to generalize. For example, the population QTE (Koenker, 2005) is a special case of the WQTE for $g(x)=1$. As another example, if $g(x)$ equals the propensity score $e(x)=\textrm{pr}(Z=1\mid X=x)$, the estimand in equation (2) corresponds to the $\tau$th QTE among the treated, a QTE analogue to the average treatment effect among the treated. See Sun et al. (2021) for more details.
@@ 2.3 Identification and estimation given complete data
$0=E\left(g(X)\frac{\mathbbm{1}\left\{Z=z\right\}}{e(X)^{z}(1-e(X))^{1-z}}\left[\mathbbm{1}\{Y\leq Q_{Y(z)}^{g}(\tau)\}-\tau\right]\right).$
These estimating equations motivate the following estimator: given a sample of independent and identically distributed draws from the joint distribution $P_{c}$,
$\mathbbm{P}_{n}\left(W^{g}(Z,X)\begin{pmatrix}1\\[-5.0pt] Z\end{pmatrix}\left[\mathbbm{1}\{Y<\beta^{g}_{0}(\tau)+\beta^{g}(\tau)Z\}-\tau\right]\right)=0,\quad\tau\in(0,1),$ (3)
with $W^{g}(Z_{i},X_{i})=g(X_{i})Z_{i}/e(X_{i})+g(X_{i})(1-Z_{i})/\{1-e(X_{i})\}$.
@@ 3.1 Double-sampling
For settings where $Y$ is potentially missing, let $R\in\{0,1\}$ denote the corresponding observance indicator so that $R=0/1$ if $Y$ is missing/observed. In this context, MAR states, intuitively, that the probability of observing $Y$ only depends on the fully observed variables. Formally, one can write MAR as: $R\perp Y\mid Z,X$. From this formal definition, MNAR may be seen to result when missingness depends on the underlying value of the outcome, either directly or through some (as-yet unspecified) other factor(s). In the bariatric surgery context, for example, the missingness process may be MNAR if patients who do not respond to surgery (i.e., fail to lose weight and/or maintain weight loss) are systematically more likely to have complete data—even after accounting for observed covariates and treatment status—perhaps because they are more likely to continue engaging with the health care system.
In cases where MAR is uncertain, we consider collecting the otherwise missing data on a sub-sample, a process termed ‘double-sampling’ by Frangakis and Rubin (2001). Let $S\in\{0,1\}$ be an indicator of double-sampling, indicating whether an individual with initially missing outcome data is selected for follow-up and complete data are collected. By design, $S=1$ only if $R=0$ and $S\equiv S(1-R)$. We define the double-sampling probability $\eta(z,x):=\textrm{pr}(S=1\mid Z=z,X=x,R=0)$ for $z\in\mathcal{Z},x\in\mathcal{X}$. Additionally, the process by which these otherwise missing data are ascertained will be context specific and may involve active participation of study subjects (e.g., response to a survey or phone call) or no participation (e.g., manual chart review); see Levis et al. (2023) for additional details. Under this scheme, the ‘observed data’ are $n$ i.i.d. copies of $\{R,S,(R+S)Y,Z,X\}$ with corresponding joint distribution $P_{o}$.
@@ 3.2 Assumptions
Here, we outline a set of assumptions used to establish identification of the effects of interest and the asymptotic properties of the estimator proposed in Section 3.3. Prior to doing so, we define $D_{1}(\tau)=E\left[g(X)f_{Y(1)\mid X}\{\beta^{g*}_{0}(\tau)+\beta^{g*}(\tau)\}\right]$ and $D_{0}(\tau)=E\left[g(X)f_{Y(0)\mid X}\{\beta^{g*}_{0}(\tau)\}\right]$, where $\tau\in(0,1)$ and $\{\beta^{g*}_{0}(\tau),\beta^{g*}(\tau)\}$ denote the true values of $\{\beta^{g}_{0}(\tau),\beta^{g}(\tau)\}$.
{assumption}
@@
{assumption}
Positivity of double-sampling probabilities: For some $c_{\eta}>0$, $c_{\eta}<\eta(z,x)<1-c_{\eta}$ for $z\in\mathcal{Z},x\in\mathcal{X}$.
{assumption}
For all $\tau\in(0,1)$, $z\in\mathcal{Z}$, the distribution function $F_{Y\mid Z=z,X}$ is differentiable at $\beta^{g*}_{0}(\tau)+\beta^{g*}(\tau)z$ with density $f_{Y\mid Z=z,X}$.
{assumption}
The derivative $y\mapsto f_{Y\mid Z,X}(y)$ is continuous and bounded in absolute value from above uniformly over $y\in\mathcal{Y}_{z,x}$ and $X,Z$.
{assumption}
@@
Assumptions 3.2, 3.2 relate directly to the double-sampling scheme, ensuring the identification of the full data distribution from the observed data. Specifically, Assumption 3.2 states that whether or not an individual is double-sampled is independent of the initially unobserved outcome, given the initially observed data. Assumption 3.2 requires that among individuals with initially missing data, the conditional probability of being double-sampled cannot be 0 or 1 in any subgroup defined by $(Z,X)$. Note, Assumption 3.2 can be viewed as an MAR assumption for the second phase of data collection. In practice, depending on the second phase sampling method, discrepancies may arise between intended sampling probabilities controlled by researchers and actual data collection methods (e.g., does an individual fill out the survey?). Thus the follow-up sample may or may not have complete data as planned. For example, in a chart review without individuals’ direct engagement, it is reasonable to anticipate that those sub-sampled will provide complete data, supporting the plausibility of Assumption 3.2. In other settings, however, such as telephone surveys, some initially selected individuals may decline to participate, leading to an additional layer of non-response (Koffman et al., 2021). In such cases, the double-sampling probabilities will have to be estimated, and the plausibility of Assumption 3.2 considered in light of this additional missingness.
@@
Assumptions 3.2, 3.2 are standard in quantile-based analyzes (Koenker, 2005) to control the asymptotic behaviour of the proposed estimator. Assumption 3.2 ensures that the expectation of the estimating equation $\phi_{\beta}$ in (7) is differentiable, allowing the use of the mean value theorem in the proof. Notably, Assumption 3.2 can be relaxed to one-sided derivatives, enabling the application of a one-sided version of the Mean Value Theorem. Assumption 3.2 requires that the conditional density function be bounded ensuring finite variance and the use of the mean value theorem. Assumption 3.2 clarifies parametric specifications for the $\eta(Z,X)$ and $e(X)$ when require estimation, acting as regularity conditions for the estimating functions $\phi_{\kappa}$ and $\phi_{\gamma}$ in Section 3.3. Implicit to the assumption is that the class of models adopted for each is sufficiently broad as to contain the true underlying mechanism. Practically, consideration of this assumption will be made in the context of whether and how flexibility is introduced into models specifications, such as using splines. While, arguably, a focus on parametric models may be restrictive, it likely aligns with how most related clinical and public health research is conducted, and so represents an important case to consider. Nevertheless, possible extensions to nonparametric models are discussed in Section 6.